OpenStudy (anonymous):

write the expression in standard form. (2+i)/(2-i)

OpenStudy (asnaseer):

ok, firstly do you know what they mean by "" here?

OpenStudy (asnaseer):

"standard form"

OpenStudy (anonymous):

a+bi right?

OpenStudy (asnaseer):

correct

OpenStudy (asnaseer):

so now think about what you can multiply the denominator of this fraction by in order to get rid of the complex part (i.e. make it a real number)

OpenStudy (anonymous):

2-i? i dont get how to clear the denominator

OpenStudy (asnaseer):

if you have some complex number of the form a + bi, then do you know what you can multiply it by in order to turn it into a real number?

OpenStudy (anonymous):

by the recipricol?

OpenStudy (asnaseer):

no. have you heard about complex conjugate?

OpenStudy (anonymous):

yes. its like z=a+bi lol thts the complex conjugate of the complex number

OpenStudy (asnaseer):

not quite. if you have a complex number of the form a + bi, then its complex conjugate will be a - bi. one of the properties of complex conjugates is that if you multiply them together you get a real number like this:\[(a+bi)(a-bi)=a^2-b^2i^2=a^2+b^2\]

OpenStudy (asnaseer):

does that look familiar at all?

OpenStudy (anonymous):

yea that does

OpenStudy (asnaseer):

ok, good. so now what would be the complex conjugate of the denominator of your fraction?

OpenStudy (asnaseer):

i.e. what is the complex conjugate of 2 - i?

OpenStudy (anonymous):

2^2 + 1^2? or is it i ^2 for b?

OpenStudy (asnaseer):

not quite, remember for any complex number of the form a + bi, its complex conjugate would be a - bi. i.e. you just change the sign of the complex part of the number.

OpenStudy (anonymous):

so 2+i?

OpenStudy (asnaseer):

perfect!

OpenStudy (asnaseer):

now remember that any number multiplied by 1 will give you the same number. so we just need to multiply your fraction by:\[\frac{2+i}{2+i}\]as this will get rid of the complex parts in the denominator. hope that makes sense?

OpenStudy (anonymous):

so you change it to the congjuate first?(the denominator) and them cancel out?

OpenStudy (asnaseer):

no no. you do this:\[\frac{2+i}{2-i}=\frac{2+i}{2-i}\times1=\frac{2+i}{2-i}\times\frac{2+i}{2+i}=\frac{(2+i)(2+i)}{(2-i)(2+i)}\]so you end up with a complex number multiplied by its conjugate in the denominator - which effectively gets rid of the complex numbers in the denominator.

OpenStudy (asnaseer):

now multiply out the numerator and the denominator

OpenStudy (anonymous):

oh ok thats sorta what i was saying though lol

OpenStudy (asnaseer):

:)

OpenStudy (anonymous):

so its 3+4i/3+4i. so its 1?

OpenStudy (asnaseer):

no - I think you made a mistake in multiplying out the denominator

OpenStudy (asnaseer):

remember that a complex number multiplied by its conjugate will always give you a real number

OpenStudy (anonymous):

denominator is (2-i )(2+i)

OpenStudy (asnaseer):

yes

OpenStudy (anonymous):

and you distribute. so 4+2i-2i-1

OpenStudy (anonymous):

so 3 ok thank you!

OpenStudy (asnaseer):

not quite

OpenStudy (asnaseer):

you got the last part wrong

OpenStudy (anonymous):

ugh!!! i suck at this

OpenStudy (asnaseer):

the last expansion should give you \(-i^2=+1\)

OpenStudy (asnaseer):

don't worry - with practice you'll soon master this :)

OpenStudy (anonymous):

oh so if its negative i it turns into +1

OpenStudy (asnaseer):

negative \(i^2\) - yes turns into plus 1 since:\[-i^2=-(i^2)=-(-1)=1\]

OpenStudy (anonymous):

so its a neagitve 3? not a positive?

OpenStudy (asnaseer):

?

OpenStudy (asnaseer):

\[(2-i)(2+i)=4+2i-2i-i^2=4-i^2=4+1=?\]

OpenStudy (anonymous):

hahah 5 nevermind :)

OpenStudy (asnaseer):

good, so the final answer in standard form will be?

OpenStudy (anonymous):

ummmm (3+4i)/5 but you need to get rid of the 5 right?

OpenStudy (asnaseer):

you need to write it in the form a + bi, so distribute the division by 5 into each of the terms in the numerator

OpenStudy (anonymous):

so 15+20i?

OpenStudy (asnaseer):

you have multiplied instead of divided

OpenStudy (asnaseer):

\[\frac{a+bi}{c}=\frac{a}{c}+\frac{b}{c}i\]

OpenStudy (anonymous):

why would you divide the 5? youre trying to get rid of denominator

OpenStudy (asnaseer):

you ended up with (3+4i)/5 - which is the right answer but is not written in standard form

OpenStudy (anonymous):

oh so what you wrote is standard form?

OpenStudy (asnaseer):

yes

OpenStudy (asnaseer):

remember when we have a + bi, a and b CAN be fractions

OpenStudy (anonymous):

ok thank you so much. sorry about the trouble

OpenStudy (asnaseer):

no trouble at all my friend - always glad to help :)

OpenStudy (asnaseer):

so - just to sure - what do you think is the final answer to this question?

OpenStudy (anonymous):

3/5 +4i/5

OpenStudy (asnaseer):

perfect!

OpenStudy (anonymous):

thanks!!

OpenStudy (asnaseer):

yw :)